# David and Goliath: An Empirical Evaluation of Attacks and Defenses for QNNs at the Deep Edge

### v1

@@ 5.1 Related Work
[17] pioneered in the evaluation of adversarial example transferability. The authors concluded that QNNs operating in int-8 or lower bit widths show defensive behavior. The authors also envision that clipping activations change the feature space of ANNs, which could offer marginal protection against transferability.
Along the same research line, i.e., adversary example transferability, [19] concluded that when the magnitude of the perturbation is small, quantization increases adversarial robustness by eliminating small perturbations; however, when the magnitude of the perturbation is greater than a given threshold, quantization acts as a distortion amplifier (quantization-shift).
[7] were the first to evaluate QNNs robustness on a non-white-box scenario (using gray-box attacks). The authors concluded that QNNs offer only marginal protection as gray-box attacks can still fool the network. Additionally, the authors state that adversarial transferability is sensitive to the quantization-shift phenomenon and the gradient misalignment.
[8] considered the findings of previous works as a baseline and defined a geometric model to evaluate QNN robustness. Through this geometric model, the authors mathematically proved that the mean distance to the decision boundary decreases as the weight precision and the input dimensionality increase. Additionally, the authors introduced the concept of critical attack strength - the attack strength at which the adversarial accuracy of QNNs matches the adversarial accuracy of ANNs.
[18] followed a different approach from previous works and proposed a training algorithm to tackle quantization-shift. More specifically, the authors proposed a quantization-aware training algorithm that maintains the Lipschitz constant of the network below 1 - the Lipschitz constant measures how much the output changes when the input changes. Additionally, the authors demonstrate that this mechanism can be combined with other defenses, such as Feature Squeezing or Adversarial Training, to enhance the robustness of QNNs.
[50] evaluated the effect of adversarial examples on 8-bit inferior QNNs. They observed that for precisions under 5 bits, adversarially trained QNNs are more vulnerable to adversarial examples than the equivalent ANN. To tackle this issue, they proposed an iteration of PGD Adversarial Training specialized for QNNs.
@@ 5.2 Gap Analysis
Attacks. Information on QNN’s adversarial robustness is scattered across multiple works that do not share the same quantization policy. In some cases, this has led to contradictory conclusions [18, 19]. While [18] states that QNNs are more vulnerable to adversarial examples, [19] suggests that QNNs show increased robustness when compared to ANNs. Nevertheless, both of these works only consider the quantization-shift phenomenon (T3) and let out all the other factors that also influence the robustness of a QNN. If we take a closer look at Table XIII, we observe that the most complete work till date comes from [7]; however, the authors do not evaluate the point distance to the decision boundary (T1) and do not follow a state-of-the-art quantization policy. Given this background, our work systematizes and empirically evaluates all the conclusions drawn from previous works, under a common quantization policy, which is widely used in TinyML (TF Lite Micro).
Compared to previous research, our empirical evaluation relies on a far more extensive list of attacks. To the best of our knowledge, we are the only ones including a parameter-free attack (AutoAttack), which reduces the chance of bad results due to human error. In the black-box scenario, which is the most realistic, we are the first to include powerful attacks. Previous works have either neglected this scenario or have only included simple attacks that add random noise to images [18, 50].
Defenses. The works [18, 50] are the only ones addressing the topic of defenses against adversarial examples for QNNs; however, in a different approach. [18] proposes a novel training algorithm to reduce the effect of quantization-shift on QNNs and couples it with adversarial training and Feature Squeezing. [50] proposes an iteration of PGD Adversarial Training targeting weight-quantized models (QNNs whose weights are quantized but activations remain in floating-point) with precision equal to or inferior to 8-bits.
@@
Although the work from [19] uses the same quantization policy as our work, the authors only evaluate the adversarial transferability vs. quantization-shift phenomenon (T3). To the best of our knowledge, our work pioneers research on adversarial example robustness for TinyML applications. Our evaluation includes three QNNs, two composing the image applications from the MLPerf Tiny benchmark [20], and another specifically developed to be executed on Arm Cortex-M MCUs [27].
@@ 6 Discussion
Although the effect of adversarial examples on QNNs is a topic that has emerged in the literature over the past few years, the information and findings available are quite scattered and sometimes contradictory at first sight. While some works claim that quantization increases the robustness against adversarial examples [19], others claim the opposite [18]. What sometimes these works fail to communicate is the type of devices they target. QNNs targeting GPUs or TPUs usually follow different quantization policies than those targeting MCUs. Our work empirically evaluates all the conclusions drawn from previous works under a common quantization policy, tailored for the TinyML scenario (TF Lite Micro). Furthermore, we are the first to empirically evaluate how current defenses tailored for full-precision ANNs perform when applied to QNNs and adapted to the TinyML environment. Our work unveils findings supported by existing literature (T1, T2, T4), extends existing knowledge (T3), and introduces entirely novel contributions (T5-T10). Although our work targets TinyML applications, only the takeaway T8 is dependent on the architecture of the target processor class. The remaining takeaways are independent of the architecture of the target processor class, i.e., they hold for any ML system following the quantization policy of TF Lite Micro. As a final remark, we envision that developing train-based defenses that can smooth the quantization-shift and gradient misalignment issues is the way to enhance the robustness of QNNs in TinyML applications.

### v2

@@ 5.1 Related Work
[…]
zhao2018 pioneered in the evaluation of adversarial example transferability. The authors concluded that QNNs operating in int-8 or lower bit widths show defensive behavior. The authors also envision that clipping activations change the feature space of ANNs, which could offer marginal protection against transferability.
[…]
bernhard2019 were the first to evaluate QNNs robustness on a non-white-box scenario (using gray-box attacks). The authors concluded that QNNs offer only marginal protection as gray-box attacks can still fool the network. Additionally, the authors state that adversarial transferability is sensitive to the quantization-shift phenomenon and the gradient misalignment.
[…]
gorsline2021 considered the findings of previous works as a baseline and defined a geometric model to evaluate QNN robustness. Through this geometric model, the authors mathematically proved that the mean distance to the decision boundary decreases as the weight precision and the input dimensionality increase. Additionally, the authors introduced the concept of critical attack strength - the attack strength at which the adversarial accuracy of QNNs matches the adversarial accuracy of ANNs.
[…]
lin2019 followed a different approach from previous works and proposed a training algorithm to tackle quantization-shift. More specifically, the authors proposed a quantization-aware training algorithm that maintains the Lipschitz constant of the network below 1 - the Lipschitz constant measures how much the output changes when the input changes. Additionally, the authors demonstrate that this mechanism can be combined with other defenses, such as Feature Squeezing or Adversarial Training, to enhance the robustness of QNNs.
[…]
song2021 evaluated the effect of adversarial examples on 8-bit inferior QNNs. They observed that for precisions under 5 bits, adversarially trained QNNs are more vulnerable to adversarial examples than the equivalent ANN. To tackle this issue, they proposed an iteration of PGD Adversarial Training specialized for QNNs.
@@ 5.2 Gap Analysis
[…]
Compared to previous research, our empirical evaluation relies on a far more extensive list of attacks. To the best of our knowledge, we are the only ones including a parameter-free attack (AutoAttack), which reduces the chance of bad results due to human error. In the black-box scenario, which is the most realistic, we are the first to include powerful attacks. Previous works have either neglected this scenario or have only included simple attacks that add random noise to images [lin2019, song2021].
[…]
@@
[…]
@@ 6 Discussion
Although the effect of adversarial examples on QNNs is a topic that has emerged in the literature over the past few years, the information and findings available are quite scattered and sometimes contradictory at first sight. While some works claim that quantization increases the robustness against adversarial examples [duncan2020], others claim the opposite [lin2019]. What sometimes these works fail to communicate is the type of devices they target. QNNs targeting GPUs or TPUs usually follow different quantization policies than those targeting MCUs. Our work empirically evaluates all the conclusions drawn from previous works under a common quantization policy, tailored for the TinyML scenario (TF Lite Micro). Furthermore, we are the first to empirically evaluate how current defenses tailored for full-precision ANNs perform when applied to QNNs and adapted to the TinyML environment. Our work unveils findings supported by existing literature (T1, T2, T4), extends existing knowledge (T3), and introduces entirely novel contributions (T5-T10). Although our work targets TinyML applications, only the takeaway T8 is dependent on the architecture of the target processor class. The remaining takeaways are independent of the architecture of the target processor class, i.e., they hold for any ML system following the quantization policy of TF Lite Micro. As a final remark, we envision that developing train-based defenses that can smooth the quantization-shift and gradient misalignment issues is the way to enhance the robustness of QNNs in TinyML applications.